Dino has 20 ounces of 20% of salt solution, how much salt should he add to make it a 25% solution
Answer by TimothyLamb(4379)

---
20*20 + 100x = 25(x + 20)
400 + 100x = 25x + 500
75x = 100
x = 100/75
---
answer:
add 1.333333 oz of pure salt to the existing 20% solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php